I hope someone can help... I have submitted 2 questions in the past any no one has replied.
my questions is;
Find the values of X and Y that solve the following system of equations:
2x-9y=-21
-3x-7y=11
So far this is what i have done;
-6x-27y=63
6x-14y=23
13y/13 = 41/13
y=41/13
2x-9y=-21
2x-9(41/13)=-21
...now what??? help, not sure how to solve with fraction

Answer by palanisamy(496)

The given equations are
2x-9y=-21 ...(1)
-3x-7y=11 ...(2)
(1)*3 => 6x-27y = -63 .. (3)
(2)*2 => -6x-14y = 22 ...(4)
(3)+(4) =? -41y= -41
y= (-41)/(-41)
y=1
(1) => 2x-9*1 = -21
2x = -21+9
2x = -12
2x= -12
x = -12/2
x = -6
The solution is (-6,1)